-3-4(t-5)=-2(t+3)+11
0.08x+0.5(x+100)=73.2

Answer by checkley77(12569)

-3-4(t-5)=-2(t+3)+11
-3-4T+20=-2T-6+11
-4T+2T=-6+11+3-20
-2T=-12
T=-12/-2
T=6 ANS.
PROOF:
-3-4(6-5)=-2(6+3)+11
-3-24+20=-2*9+11
-7=-18+11
-7=-7
-----------------------
0.08x+0.5(x+100)=73.2
.08X+.05X+50=73.2
..58X=73.2-50
.58X=23.2
X=23.2/.58
X=40 ANS.
PROOF:
.08*40+.05(40+100)=73.2
3.2+.5*140=73.2
3.2+70=73.2
73.2=73.2